# Combining Serverless and High-Performance Computing Paradigms to support ML Data-Intensive Applications

### v2

@@ Abstract
Data is found everywhere, from health and human infrastructure to the surge of sensors and the proliferation of internet-connected devices. To meet this challenge, the data engineering field has expanded significantly in recent years in both research and industry. Traditionally, data engineering, Machine Learning, and AI workloads have been run on large clusters within data center environments, requiring substantial investment in hardware and maintenance. With the rise of the public cloud, it is now possible to run large applications across nodes without owning or maintaining hardware. Serverless functions such as AWS Lambda provide horizontal scaling and precise billing without the hassle of managing traditional cloud infrastructure. However, when processing large datasets, users often rely on external storage options that are significantly slower than direct communication typical of HPC clusters. We introduce Cylon, a high-performance distributed data frame solution that has shown promising results for data processing using Python. We describe how we took inspiration from the FMI library and designed a serverless communicator to tackle communication and performance issues associated with serverless functions. With our design, we demonstrate that the performance of AWS Lambda falls below one percent of strong scaling experiments compared to serverful AWS (EC2) and HPCs based on implementing direct communication via NAT Traversal TCP Hole Punching.
@@ I Introduction and Motivation
In this paper, we will explore our high-performance serverless, a runtime engine designed to optimize the execution of deep learning applications in high-performance computing (HPC) and cloud environments. We introduce a serverless communicator design that leverages NAT Hole Punching to provide a high-performance alternative to intermediate storage solutions like AWS S3. This enables function-to-function bidirectional communication, a capability not natively supported by AWS Lambda. This is also an ideal choice for AWS Fargate, serverless Docker on AWS ECS, based on the difficulties in implementing high-performance communication using libraries such as UCX, MPI, and Gloo. We introduce and incorporate Cylon to achieve an existing goal of parallel computing, making it achievable without user intervention. This is accomplished by utilizing libraries of previously parallelized operators (such as join and all the other Pandas operators, as well as deep learning). We will explain the advantages of executing workloads within Docker containers on AWS using ECS tasks and discuss how this approach can be extended to HPC environments utilizing Singularity or Apptainer so that it is possible to execute derivative experiments.
@@
Next, we will address the challenges of large-scale parallel execution on AWS by integrating direct communication support via NAT Traversal TCP Hole Punching in Cylon for pre-processing and inference in serverless environments. By demonstrating the effectiveness of our approach, we will highlight its applicability in scientific fields like bioinformatics, hydrology, earthquake prediction, astronomy, and genome research, where it can address high-performance communication issues observed in current state-of-the-art approaches, such as distributed storage latency and data parallelism. Furthermore, we will explore the cost savings associated with FaaS execution to showcase the relevance and innovation of this approach compared to serverful execution in traditional HPC environments. Our experimental results demonstrate that our serverless communicator design achieves performance that approximates EC2 with deviations of less than 1% for both weak and strong scaling experiments, significantly outperforming traditional object storage approaches such as AWS S3 used in current state-of-the-art serverless bioinformatics applications.
@@ III-C Serverless, High-Performance Architecture Using Cylon
As part of this work, we have designed a high-performance serverless architecture within Cylon based on a reference FMI interface to support the direct communication necessary for Bulk Synchronous Parallel architectures. As part of this work, We will discuss a reference serverful architecture for running data-parallel applications on HPCs and traditional cloud services such as EC2 using Cylon, and outline our design that introduces a high-performance serverless communication that can be used with serverless infrastructure such as AWS Fargate and AWS Lambda.
@@ III-E Serverless High-Performance BSP Architecture
Figure 6 illustrates our architecture. Our design calls for the use of an AWS Step Function to encapsulate the AWS Serverless Execution Engine, as shown in Figure 7. AWS Step Functions are a Software as a Service (SaaS) feature that facilitates the orchestration of AWS services into serverless workflows. The Lambda: init function receives a payload containing various parameters, such as the number of rows, world size, number of iterations, target S3 bucket, script, and output paths. We utilize a specific S3 bucket to store invocation results in a folder using the Boto3 library. This approach is advantageous over parsing the data output from log files in AWS CloudWatch, a SaaS metrics repository created and utilized by AWS services, as it would be cumbersome. The Lambda invoke state forwards this input JSON payload to a cylon_init AWS Lambda function. This function validates the availability of the rendezvous server. It generates an N-sized array of JSON payloads, which is then sent to the Map, ExtractAndInvokeLambda states, which is configured to use distributed processing mode to enable highly parallel execution. Next, each generated payload is forwarded to the Lambda: Invoke state from the pass ExtractAndInvokeLambda state. This state passes the input to output, executing the fmi_executor AWS Lambda function. The runtime of the cylon_executor function includes the high-performance MLSys runtime described earlier, which consists of a communicator that supports direct communication over a NAT gateway described earlier. The incoming payload specifies the location of the target script in the S3 bucket. Upon execution of the Lambda function, the contents of the specified S3 bucket are retrieved and persisted in temporary ephemeral storage. Subsequently, Python is used to execute the script used for our experiments. Our design supports both a single script and a folder containing Python scripts and modules.
@@ IV-A Join Operation Scalability
S is the Speedup: S = $\frac{T_{1}}{T_{2}}$
$\Delta{S}$ is the error in speedup
$T_{1}$ is the baseline execution time (the time for one node)
$\Delta{T}_{1}$ is the error in baseline execution time
$T_{2}$ is the parallel execution time
$\Delta{T}_{2}$ is the error in parallel execution time
Figure 8 depicts the weak scaling experiments. In an ideal scenario, we should see close to constant performance. However, based on system overhead, we mitigate variation by executing ten iterations for each experiment. In general, as parallelism is increased, resource and scheduling overhead contribute to the decreasing performance across weak scaling experiments. We observe that our serverless communicator design resulted in performance that approximates EC2 weak scaling with deviation(s) of less than 1 percent.
@@
Table III presents the results of strong scaling experiments involving 4.5 million rows. Like weak scaling, we conducted each experiment four times, performing ten iterations for each trial. The results of strong scaling are illustrated in Figure 9. We observe high latencies across the infrastructure for fewer than four nodes. However, we also observe a general convergence of latency as the number of nodes ranges from 16 to 64. This supports utilizing cloud resources as a viable alternative for data-parallel pre-processing workloads. We observe that the performance of AWS Lambda falls below one percent of strong scaling experiments with EC2 and Rivanna, based on implementing direct communication via NAT Traversal TCP Hole Punching for serverless architectures such as AWS Lambda and AWS Fargate.
@@ VI Conclusion
In this paper, we detail our design of UCX and UCC as an alternative to MPI in BSP communication between processes based on the clear applicability to serverful cloud and HPC environments. The novelty of this work is apparent based on a survey of the literature on similar work involving distributed data frames. We applied this work directly and refined the initial implementation to support both serverful and serverless architectures of cloud providers, such as Amazon Web Services. Additionally, we implemented direct support for Python via Cython, enabling the execution of data engineering tasks in Python without the need to write native code. We also completed the integration of UCC into Cylon by removing our implementation of AllGatherV in favor of UCC’s support for variable-length data. Furthermore, we included container support across infrastructures to facilitate the execution of experiments and library usage. On the serverless side, we detailed our serverless communicator, drawing inspiration from the FMI library, and conducted experiments to compare this design with our serverful design.
During this research, we encountered several challenges. We first validated the execution of Cylon on AWS by building the library using AWS Workspaces, which was generally successful. Using AWS facilitated the building and running on bare metal due to complete control of the underlying ECS host. Comparatively, building on bare metal on HPCs was problematic based on the library versions supported. We solved this problem by using Apptainer or Singularity and running Cylon in containers on HPC systems.
@@
Second, the reference FMI library did not include an All-To-All collective operation that we needed for Cylon communicator integration. Second, it was necessary to include variable allgather, variable gather, support for non-blocking operations, retries for socket connection failures, a ping/ping capability to prevent eager termination of sockets, and rank calculation using atomic counters. An initial challenge was coordination between nodes and associated race conditions. Rank-based ordering for blocking operations using Redis locks was incorporated into our design to address race conditions observed during blocking socket operations. Note that FMI, in its original form, only supports blocking operations, and our design called for non-blocking I/O. These additions resulted in improvements and a robust communicator that can be leveraged across serverless platforms.
In this paper, we demonstrate experiments using Cylon and UCC/UCX, comparing them to similar experiments conducted on the University of Virginia’s Rivanna HPC cluster to validate the effectiveness and applicability of running data engineering tasks at scale on clouds. We discussed the challenges of using serverless technology, particularly in terms of communication and coordination. We describe and design an approach that utilizes NAT traversal through NAT hole punching to enable direct communication between AWS Lambda functions. Additionally, we highlight the relevance of this work to serverless containers and emphasize the importance of architecting communication layers at high levels of abstraction when executing ML workloads on serverless compute. We demonstrated the effectiveness of NAT Traversal through NAT hole punching and observed strong scaling performance similar to what is achievable on HPCs such as Rivanna in Cylon. We believe that this approach can be applied to bioinformatics analysis, genome research, hydrology, astronomy, earthquake prediction, and other time series-based foundational models, as surveyed in the current state of the art in these fields.
@@ Acknowledgments
We gratefully acknowledge the support of the Department of Energy and the National Science Foundation through DE-SC0023452, NSF-2421782, NSF - 2504401, and NSF-2411009 grants.

### v3

@@ Abstract
Data is found everywhere, from health and human infrastructure to the surge of sensors and the proliferation of internet-connected devices. To meet this challenge, the data engineering field has expanded significantly in recent years in both research and industry. Traditionally, data engineering, Machine Learning, and AI workloads have been run on large clusters within data center environments, requiring substantial investment in hardware and maintenance. With the rise of the public cloud, it is now possible to run large applications across nodes without owning or maintaining hardware. Serverless functions such as AWS Lambda provide horizontal scaling and precise billing without the hassle of managing traditional cloud infrastructure. However, when processing large datasets, users often rely on external storage options that are significantly slower than direct communication typical of HPC clusters. We introduce Cylon, a high-performance distributed data frame solution that has shown promising results for data processing using Python. We describe how we took inspiration from the FMI library and designed a serverless communicator to tackle communication and performance issues associated with serverless functions. With our design, we demonstrate that the scaling efficiency of AWS Lambda achieves within 6.5% of serverful AWS (EC2) at 64 nodes, based on implementing direct communication via NAT Traversal TCP Hole Punching.
@@ I Introduction and Motivation
In this paper, we will explore our high-performance serverless, a runtime engine designed to optimize the execution of deep learning applications in high-performance computing (HPC) and cloud environments. We introduce a serverless communicator design that leverages NAT Hole Punching to provide a high-performance alternative to intermediate storage solutions like AWS S3. This enables function-to-function bidirectional communication, a capability not natively supported by AWS Lambda. This is also an ideal choice for AWS Fargate, serverless Docker on AWS ECS, based on the difficulties in implementing high-performance communication using libraries such as UCX, MPI, and Gloo. We introduce and incorporate Cylon to achieve an existing goal of parallel computing, making it achievable without user intervention. This is accomplished by utilizing libraries of previously parallelized operators (such as join and other Pandas operators, as well as deep learning). We will explain the advantages of executing workloads within Docker containers on AWS using ECS tasks and discuss how this approach can be extended to HPC environments utilizing Singularity or Apptainer so that it is possible to execute derivative experiments.
@@
Next, we will address the challenges of large-scale parallel execution on AWS by integrating direct communication support via NAT Traversal TCP Hole Punching in Cylon for pre-processing and inference in serverless environments. By demonstrating the effectiveness of our approach, we will highlight its applicability in scientific fields like bioinformatics, hydrology, earthquake prediction, astronomy, and genome research, where it can address high-performance communication issues observed in current state-of-the-art approaches, such as distributed storage latency and data parallelism. Furthermore, we will explore the cost savings associated with FaaS execution to showcase the relevance and innovation of this approach compared to serverful execution in traditional HPC environments. Our experimental results demonstrate that our serverless communicator design achieves scaling efficiency within 6.5% of EC2 at 64 nodes for strong scaling experiments, significantly outperforming traditional object storage approaches such as AWS S3 used in current state-of-the-art serverless bioinformatics applications.
I-A Contributions
This work makes the following research contributions:
(i) BSP Execution Model for Serverless Computing: We demonstrate that Bulk Synchronous Parallel (BSP) workloads, traditionally requiring dedicated high-performance computing (HPC) infrastructure with low-latency interconnects, can effectively execute in serverless environments. By enabling direct peer-to-peer communication between Lambda functions through NAT traversal, we demonstrate that parallel computations achieve scaling efficiency within 6.5% of that observed on provisioned cloud virtual machines (VMs). This challenges the assumption that serverless computing is limited to embarrassingly parallel or loosely coupled workloads.
(ii) Unified Framework for ML Data Pipelines Across Computing Paradigms: We introduce Cylon, a portable and distributed dataframe library that allows the same data-intensive operations to run seamlessly across various environments, including serverless (AWS Lambda), cloud virtual machines (EC2), and high-performance computing (Rivanna). This capability supports both machine learning training and inference pipelines.
Scientific fields with significant data processing requirements, such as genomics, hydrology, and astronomy, can benefit from Cylon’s distributed operations in serverless deployments. This enables elastic scaling for fluctuating workloads without the need for dedicated infrastructure.
(iii) Quantitative Cost-Performance Analysis for Serverless HPC: We present an empirical cost model for data-intensive serverless workloads. Our analysis shows that serverless can be cost-competitive for bursty workloads: a 32-worker distributed join costs approximately \$0.03 on Lambda with pay-per-millisecond billing, compared with provisioned EC2 instances where idle time dominates cost for intermittent workloads.
[…]
@@ III-C Serverless, High-Performance Architecture Using Cylon
As part of this work, we have designed a high-performance serverless architecture within Cylon based on a reference FMI interface to support the direct communication necessary for Bulk Synchronous Parallel architectures. As part of this work, We will discuss a reference serverful architecture for running data-parallel applications on HPCs and traditional cloud services such as EC2 using Cylon, and outline our design that introduces a high-performance serverless communicator that can be used with serverless infrastructure such as AWS Fargate and AWS Lambda.
@@ III-E Serverless High-Performance BSP Architecture
Figure 6 illustrates our architecture. Our design calls for the use of an AWS Step Function to encapsulate the AWS Serverless Execution Engine, as shown in Figure 7. AWS Step Functions are a Software as a Service (SaaS) feature that facilitates the orchestration of AWS services into serverless workflows. The Lambda: init function receives a payload containing various parameters, such as the number of rows, world size, number of iterations, target S3 bucket, script, and output paths. We utilize a specific S3 bucket to store invocation results in a folder using the Boto3 library. This approach is advantageous over parsing the data output from log files in AWS CloudWatch, a SaaS metrics repository created and utilized by AWS services, as it would be cumbersome. The Lambda invoke state forwards this input JSON payload to a cylon_init AWS Lambda function. This function validates the availability of the rendezvous server. It generates an N-sized array of JSON payloads, which is then sent to the Map, ExtractAndInvokeLambda states, which is configured to use distributed processing mode to enable highly parallel execution. Next, each generated payload is forwarded to the Lambda: Invoke state from the pass ExtractAndInvokeLambda state. This state passes the input to output, executing the cylon_executor AWS Lambda function. The runtime of the cylon_executor function includes the high-performance MLSys runtime described earlier, which consists of a communicator that supports direct communication over a NAT gateway described earlier. The incoming payload specifies the location of the target script in the S3 bucket. Upon execution of the Lambda function, the contents of the specified S3 bucket are retrieved and persisted in temporary ephemeral storage. Subsequently, Python is used to execute the script used for our experiments. Our design supports both a single script and a folder containing Python scripts and modules.
@@ IV-A Join Operation Scalability
$S$ $=\text{the Speedup, }\frac{T_{1}}{T_{2}}$
$\Delta S$ $=\text{the error in speedup}$
[…]
$\Delta T_{1}$ $=\text{the error in baseline execution time}$
$T_{2}$ $=\text{Parallel execution time}$
$\Delta T_{2}$ $=\text{Error in parallel execution time}$
@@
Figure 8 depicts the weak scaling experiments. In an ideal scenario, we should see close to constant performance. However, based on system overhead, we mitigate variation by executing ten iterations for each experiment. In general, as parallelism is increased, resource and scheduling overhead contribute to the decreasing performance across weak scaling experiments. We observe that our serverless communicator design resulted in performance that closely approximates EC2 weak scaling behavior. Table III presents the results of strong scaling experiments involving 4.5 million rows. Like weak scaling, we conducted each experiment four times, performing ten iterations for each trial. The results of strong scaling are illustrated in Figure 9. We observe high latencies across the infrastructure for fewer than four nodes. However, we also observe a general convergence of latency as the number of nodes ranges from 16 to 64. This supports utilizing cloud resources as a viable alternative for data-parallel pre-processing workloads. We observe that the scaling efficiency of AWS Lambda achieves within 6.5% of EC2 at 64 nodes, based on implementing direct communication via NAT Traversal TCP Hole Punching for serverless architectures such as AWS Lambda and AWS Fargate.
@@
[…]
Table IV: Strong Scaling Speedup Comparison: Lambda vs EC2
Nodes EC2 Speedup Lambda Speedup Difference
1 $\times 1.00$$\times$ 0.0%
[…]
The faster single-node execution time observed on Rivanna compared to EC2 is attributed to compute differences: Rivanna uses Intel Xeon Gold 6248 (Cascade Lake, 2019) while EC2 m3.xlarge instances use Intel Xeon E5-2670 v2 (Ivy Bridge, 2013). The newer Cascade Lake architecture provides approximately 40% better instructions per cycle (IPC) for Arrow/join kernels. This compute difference does not affect our scaling conclusions, as we compare scaling efficiency (speedup ratios) rather than absolute performance.
IV-B Communication Infrastructure Comparison
To quantify the benefit of our direct NAT hole-punching approach, we conducted baseline comparisons against storage-mediated communication alternatives commonly used in serverless environments. Figure 10 shows weak scaling results for the Join operation using three communication channels on AWS Lambda: direct TCP via NAT traversal, Redis-mediated, and S3-mediated message passing.
Figure 10: Communication Infrastructure Comparison: Direct TCP (NAT hole-punching) vs Redis vs S3 for distributed Join on AWS Lambda
[…]
IV-C Additional Data Engineering Operators
To demonstrate that our serverless architecture supports data engineering operations beyond Join, we conducted weak scaling experiments for the GroupBy aggregation operator on AWS Lambda. GroupBy uses the same shuffle (all-to-all) communication pattern as Join, partitioning data by group key columns before performing local aggregation.
Figure 11: GroupBy weak scaling on AWS Lambda using direct TCP communication
[…]
IV-D Communication Microbenchmarks
To isolate communication overhead from application-level computation, we conducted microbenchmarks measuring the performance of MPI-style collective operations on AWS Lambda. These benchmarks are particularly relevant to distributed ML workloads, where AllReduce is fundamental to data-parallel gradient aggregation.
Figure 12: AllReduce latency vs message size on AWS Lambda
Figure 12 shows AllReduce latency across message sizes from 8 bytes to 1 MB. Latency remains relatively flat across message sizes, indicating latency-bound rather than bandwidth-bound behavior for most operations. At 32 nodes, AllReduce completes in approximately 13 milliseconds, demonstrating that MPI-style collectives can achieve millisecond-scale latency in serverless environments.
Figure 13: Barrier latency scaling on AWS Lambda
Figure 13 shows Barrier latency scaling with node count. Barrier latency scales with $\log_{2}(N)$ as expected for a binomial tree algorithm: 0.9ms at 2 nodes, 2.7ms at 8 nodes, and 7ms at 32 nodes. This confirms that synchronization overhead remains in the single-digit millisecond range, fast enough for BSP workloads requiring frequent barriers.
IV-E Serverless Execution Time Composition
Figure 14 shows the breakdown of execution time into three phases: initialization (connection setup via NAT traversal), data generation, and computation. This breakdown provides insight into where time is spent in serverless execution and helps explain performance differences across platforms.
Figure 14: Serverless execution time composition: initialization, data generation, and computation phases
At 32 nodes, the NAT traversal initialization phase dominates wall-clock time for direct TCP communication, requiring approximately 31.5 seconds for connection setup. This connection establishment time scales linearly with the number of tree levels in the binomial connection algorithm. In contrast, storage-mediated channels (Redis, S3) have negligible initialization overhead but incur higher per-operation latency during computation.
IV-F Cost Analysis
A key motivation for serverless computing is cost optimization for bursty workloads. We implemented a comprehensive cost-tracking framework to provide empirical cost analysis for our experiments, capturing Lambda compute cost (GB-seconds), Step Functions orchestration cost (state transitions), and data transfer costs.
Figure 15: Serverless execution cost analysis: Lambda compute and Step Functions orchestration
Figure 15 shows the cost breakdown across node counts and operations. Step Functions orchestration cost (shown as the lighter portion) is negligible compared to Lambda compute cost. At 32 nodes, a Join using Redis-mediated communication costs approximately \$0.032, demonstrating the cost-effectiveness of serverless for distributed data operations.
Figure 16: Cost comparison by operation and communication channel
[…]
Despite the overhead costs, serverless remains highly cost-effective for bursty, intermittent workloads where provisioned infrastructure would incur idle time charges. The total cost for all revision experiments (120 Lambda executions across 5 experiment types) was only \$3.25.
@@ VI Conclusion
In this work, we demonstrated that BSP-style data-intensive workloads can execute efficiently in serverless environments by rethinking the communication substrate. By integrating direct communication through NAT traversal and TCP hole punching into Cylon, we enabled message passing among AWS Lambda functions that avoids the latency penalties of storage-mediated approaches such as object stores or Redis. Our experimental results show that AWS Lambda achieves scaling efficiency within 6.5% of EC2 at 64 nodes, with 10–100× lower latency than storage-based communication. Our work further provides a quantitative cost-performance analysis, showing serverless to be cost-competitive for bursty workloads, and compares communication substrates (direct TCP, Redis, and S3), establishing NAT-based communication as significantly lower latency (10–100× faster) than storage-mediated alternatives. These findings challenge the prevailing assumption that serverless computing is limited to embarrassingly parallel workloads and establish that high-performance distributed data engineering and ML preprocessing are feasible in FaaS environments.
@@
We detail our design of UCX and UCC as an alternative to MPI in BSP communication between processes based on the clear applicability to serverful cloud and HPC environments. The novelty of this work is apparent based on a survey of the literature on similar work involving distributed data frames. We applied this work directly and refined the initial implementation to support both serverful and serverless architectures of cloud providers, such as AWS. Additionally, we implemented direct support for Python via Cython, enabling the execution of data engineering tasks in Python without the need to write native code. We also completed the integration of UCC into Cylon by removing our implementation of AllGatherV in favor of UCC’s support for variable-length data. Furthermore, we included container support across infrastructures to facilitate the execution of experiments and library usage. On the serverless side, we detailed our serverless communicator, drawing inspiration from the FMI library, and conducted experiments to compare this design with our serverful design.
@@
During this research, we faced several challenges. We first validated Cylon on AWS and addressed deployment issues by using Apptainer (Singularity) containers on HPC systems. Second, the reference FMI library did not include an All-To-All collective operation that we needed for Cylon communicator integration. Third, it was necessary to include variable allgather, variable gather, support for non-blocking operations, retries for socket connection failures, a ping/ping capability to prevent eager termination of sockets, and rank calculation using atomic counters. An initial challenge was coordination between nodes and associated race conditions. Rank-based ordering for blocking operations using Redis locks was incorporated into our design to address race conditions observed during blocking socket operations. Note that FMI, in its original form, only supports blocking operations, and our design called for non-blocking I/O. These additions resulted in improvements and a robust communicator that can be leveraged across serverless platforms.
Our approach can be applied to bioinformatics analysis, genome research, hydrology, astronomy, earthquake prediction, and other time series-based foundational models, as surveyed in the current state of the art in these fields.
@@ Acknowledgments
We gratefully acknowledge the support of the Department of Energy and the National Science Foundation through DE-SC0023452 FAIR Surrogate Benchmarks Supporting AI and Simulation Research, NSF-Simons AI Institute for Cosmic Origins (CosmicAI: Grant 2421782) Seed Grant, NSF-2200409 for CyberTraining:CIC: CyberTraining for Students and Technologies from Generation Z. NSF-2504401 OAC Core: RINAS: Data I/O CyberInfrastructure for Extreme-scale Foundation Model and Generative AI Training on HPC.